If the revenue function for selling q widgets is r(q) =q.(4187-5q) dollars, find the marginal revenue function, and find the marginal revenue when q=135
MR=R'(q)= ?
R'(135) =?

Answer by amarjeeth123(569)

The revenue function for selling q widgets is R(q)=q(4187-5q) dollars
Marginal revenue function=MR=R'(q)=4187-5q+q(-5)=4187-10q
Marginal revenue R'(135)=4187-10(135)=4187-1350=2837 dollars